ADA Pharmaceutical Company produces three drugs—Diomycin, Homycin, and Addolin—belonging to the analgesic (pain-killer) family of medication. Since its inception four years ago, ADA has used a direct labor hour–based system to assign manufacturing overhead costs to products.

Eme Weissman, the president of ADA Pharmaceutical, has just read about activity-based costing in a trade journal. With some curiosity and interest, she asked her financial controller, Takedo Simon, to examine differences in product costs between the firm’s current costing and activity-based costing systems.

ADA has the following budget information for the year:

Diomycin Homycin Addolin
Cost of direct materials $ 197,000 $ 257,000 $ 250,000
Cost of direct labor 242,000 226,000 255,000
Number of direct labor hours 7,120 6,720 1,920
Number of capsules 920,000 492,000 292,000

ADA has identified the following activities and cost drivers and has assigned them a total overhead cost of $196,000.

Activity Cost Driver Budgeted Overhead Cost Budgeted Cost Driver Volume
Machine setup Setup hours $ 15,200 1,520
Plant management Number of Workers 35,200 1,120
Supervision of direct labor Direct labor hours 45,200 15,200
Quality inspection Inspection hours 49,600 1,010
Order expediting Customers served 50,800 605
Total overhead $ 196,000

Takedo selected the cost drivers with the following justifications:

SETUP HOURS: The cost driver of setup hours is used because the same product takes about the same amount of setup time regardless of size of batch. For different products, however, the setup time varies.

NUMBER OF WORKERS: Plant management includes plant maintenance and corresponding managerial duties that make production possible. This activity depends on the number of workers. The more workers involved, the higher the cost.

DIRECT LABOR HOURS: Supervisors spend their time supervising production. The amount of time they spend on each product is proportional to the direct labor hours worked.

INSPECTION HOURS: Inspection involves testing a number of units in a batch. The time varies for different products but is the same for all similar products.

NUMBER OF CUSTOMERS SERVED: The need to expedite production increases as the number of customers served by the company increases. Thus, the number of customers served by ADA is a good measure of expediting production orders.

Takedo gathered the following information about the cost driver volume for each product:

Diomycin Homycin Addolin
Setup hours 160 560 760
Number of workers 160 360 560
Direct labor hours 7,120 6,720 1,920
Inspection hours 110 160 660
Customers served 37 92 460

rev: 09_27_2018_QC_CS-140513

Required:

1. Use the firm’s current costing system to calculate the unit cost of each product.

2. Use the activity-based cost system to calculate the unit cost of each product.
Answer #1

1) Absorbing overhead on the basis of of  direct labor hour

total overhead = 196,000

total labour hours

Diomycin = 7120

homycin = 6720

Addolin =1920

15760

overhead absorbtion rate(OAR)= 196,000/15760

= 12.44/ hour   material cost per unit for diomycin=197000/920000

= .214

  ( similarly can be founded for all medicine in material and labour)

cost card

   Diomycin    homycin Addolin   .0962 =(7120/920000)*12.44   

material .214 .522 .856

labour .263 .459 .873

overhead .096 .170 .082
  
cost per unit .573 1.15 1.81

2) using Activity based costing

setup cost for diomycin = (15200/1520)*160 =1600

(similarly every cost pool can be founded for each medicines)

  Diomycin    homycin Addolin   

    machine setup 1600 5600 7600

plant management 5028.8 11314.8 17600.8

Supervision of direct labor 21146.4 19958.4 5702.4

Quality inspection 5402.1 7857.6 32412.6

Order expediting 3106.9 7725.2 38626.2

total oh /drug        36284.2 52456 101942

no of unit 920000 492000 292000

OH/drug .0394 .107 .349

cost card

   Diomycin    homycin Addolin    

material .214    .522    .856

labour .263   .459   .873

overhead   .0394   .107 .349
  
cost per unit .5164 1.088 2.081